OpenStudy (anonymous):

find all the points of intersection: y=cosx and y=1- sin(x/2)

OpenStudy (anonymous):

One way to do it: \[\begin{align*} \cos x&=1-\sin\frac{x}{2}\\\\ \sin\frac{x}{2}&=1-\cos x\\\\ \left(\sin\frac{x}{2}\right)^2&=(1-\cos x)^2&\text{half-angle identity for next step}\\\\ \frac{1}{2}-\frac{1}{2}\cos x&=1-2\cos x+\cos^2x\\\\ \cos^2x-\frac{3}{2}\cos x-\frac{1}{2}&=0\\\\ 2\cos^2x-3\cos x-1&=0\end{align*}\]

OpenStudy (anonymous):

Subscribe to Whimsical Wiggy on Youtube.com to laugh at my funny stuff. I also got bacon.